OpenStudy (anonymous):

help:

OpenStudy (anonymous):

Can someone explain how to manipulate this ratio: \[\frac{F _{1}}{F_{2}} \alpha \left( \frac{r}{2r} \right)^2\] so that F _{2} = 1/4 of F _{1}. I hope that made sense.

OpenStudy (anonymous):

\[\left (\frac{r}{2r}\right )^2=\left (\frac{1}{2}\right )^2=\frac{1}{4}\]\[\frac{F_1}{F_2}~\frac{\alpha}{4}=1\]\[F_2=\alpha \frac{1}{4}F_1\]

OpenStudy (anonymous):

ahh of course, how stupid I am. Thank you!!